### v1

@@ Abstract
Context:Eccentric exoplanets offer an opportunity to study the response of an atmosphere to changing thermal forcing and the robustness of the superrotating equatorial jet seen on tidally locked hot Jupiters. However, the atmospheric dynamics on eccentric planets strongly depend on the planetary rotation period, which is difficult to constrain observationally. The ringing phenomenon, whereby the observed emission increases and decreases after the periastron passage as the flash-heated hemisphere rotates in and out of view, can provide a tight constraint on rotation.
Aims:We examine which of the highly eccentric transiting exoplanets HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b and HD 17156 b displays sufficiently strong ringing signals for the James Webb Space Telescope (JWST) to detect.
Methods:We implemented the treatment of eccentricity and non-synchronous rotation in the non-grey climate model expeRT/MITgcm and generate synthetic light curves.
Results:We find four detectable ringing peaks on HD 80606 b and some undetectable ringing on TOI-4127 b and HD 17156 b. The lack of clouds, photo-chemistry and obliquity in our models may lead us to overestimate the amplitude of the ringing however. The strength of the ringing signal is mostly determined by the eccentricity, planetary rotation period, planet-to-star radius ratio and apparent magnitude of the system. We seek more exoplanets that could show ringing but find no candidates as promising as HD 80606 b.
Conclusions:We recommend prioritising HD 80606 b as a target for ringing with JWST. A baseline of five days after the periastron passage captures three ringing peaks, sufficient to tightly constrain the planetary rotation period. An extension to seven days adds a fourth peak, verifying the rotation period.
Examining the detectability of ringing on highly eccentric exoplanets
@@
In the past few decades there have been great advances in the study of exoplanetary atmospheres. The observations of the recently launched James Webb Space Telescope (JWST) have contributed greatly to these efforts already (e.g. JWST Transiting Exoplanet Community Early Release Science Team et al., 2023; Feinstein et al., 2023; Tsai et al., 2023a; Dyrek et al., 2024). Because of the complicated nature of exoplanetary atmospheres, observations ought to be compared to global circulation models (GCMs) to be interpreted properly. These GCMs require well constrained orbital and planetary parameters, such as the planet’s radius, rotation period and orbital eccentricity, as input Showman et al. (2020). If the planet transits its parent star most of these quantities can be reliably measured with the notable exception of the rotation period (Sing, 2018).
@@
Most studies so far have focussed on hot Jupiters because their size and brightness makes it relatively easy to achieve a high observational signal-to-noise (Showman et al., 2020). As a result of their close-in orbit, hot Jupiter are often on a tidally circularised orbit. Based on timescale arguments it can then safely be assumed that the planet’s rotation period is equal to the orbital period (Hut, 1981; Socrates et al., 2012; Rauscher, 2017) and its obliquity is zero (Rauscher, 2017; Rauscher et al., 2023). Consequently, the irradiation these planets receive is constant in intensity and always illuminates the same hemisphere. On the dayside hemisphere, the illumination creates a pattern of geostrophic winds that transports angular momentum towards the equator (Showman & Polvani, 2010, 2011). Consequently, a remarkable circulation pattern forms dominated by an equatorial superrotating jet, an eastwards flow of gas centred on the equator extending around the planet and typically reaching speeds of several km s. Such jets are well established both theoretically (e.g. Showman & Guillot, 2002; Showman et al., 2009; Tsai et al., 2014; Amundsen et al., 2016; Carone et al., 2020; Showman et al., 2020) and observationally (e.g. Knutson et al., 2007; Snellen et al., 2010; Zellem et al., 2014; Louden & Wheatley, 2015; May et al., 2022). Because of the major role of the Coriolis force in atmospheric dynamics, the planet’s rotation period is a decisive parameter to these circulation patterns (e.g. Showman et al., 2009; Showman & Polvani, 2010; Rauscher & Kempton, 2014; Showman et al., 2015; Baeyens et al., 2021; May et al., 2022).
@@
In the case of a planet on an eccentric orbit, it can no longer be assumed to rotate synchronously Hut (1981). In spite of the increased difficulty, studying eccentric and non-synchronously rotating exoplanets is worthwhile as it allows one to examine the response of the atmosphere’s dynamics, chemistry and cloudiness to changing thermal forcing (Kataria et al., 2013; Lewis et al., 2017; Tsai et al., 2023b). Moreover, transit spectroscopy of cool planets on a far-out, circular orbit is rarely available as the distance restricts the viewing angles from which the planet would transit. On an eccentric planet on the other hand, a transit is more likely to take place around periastron where the atmosphere still contains traces of the cooler regime around apastron that is otherwise difficult to observe. Finally, about half of the known Jupiter-like planets are eccentric, especially those with a longer orbital period due to the distance dependence of tidal circularisation (Malmberg & Davies, 2009; Dawson & Murray-Clay, 2013; Bitsch et al., 2020).
@@
A number of previous studies have already examined the role of eccentricity and non-synchronous rotation on the atmospheric circulation of hot and warm Jupiters. Showman et al. (2009); Rauscher & Kempton (2014); Showman et al. (2015) used 3D GCMs to study the effect of non-synchronous rotation, concluding that a shorter rotation period weakens the equatorial jet. They disagree on the effect of increasing the rotation period though, as Rauscher & Kempton (2014) claim that a westward flow develops while Showman et al. (2009, 2015) find that it leads to a significantly broader and faster eastward jets. Nonetheless, these studies all show that the rotation period is of major importance to atmospheric circulation. In a similar vein Kataria et al. (2013) examined the impact of eccentricity in a theoretical scenario, pointing out that while jets are still ubiquitous they tend to be faster the greater the eccentricity is and change quite notably in the upper atmosphere over the orbit. This is in agreement with Lewis et al. (2014) who simulated the eccentric planet HAT-P-2 b and Lewis et al. (2017); Tsai et al. (2023b) who examined the extremely eccentric HD 80606 b. The latter studies show that the jet speed on HD 80606 b decreases between periastron passages. Lewis et al. (2017) also highlight the impact of the rotation period on the circulation, composition and lofting of clouds and consequently the photometric observations of highly eccentric planets.
@@
During the periastron passage when an eccentric planet is briefly but intensely heated, most of the flash-heating is deposited unto the side facing the star. As the planet moves further from the star and cools, that heated hemisphere rotates into and out of view of an observer on Earth. Consequently, the total brightness of the star and planet combined may display a number of local maxima, which is known as ringing. Thus, by studying the system’s light curve after the periastron passage, one can constrain the planet’s rotation period as well as how efficiently the temperature is homogenised by circulation. Using a simple thermal model without consistent dynamics or radiative transfer, Cowan & Agol (2011) predicted the presence of such ringing on the exoplanets HAT-P-2 b, HD 80606 b and HD 17156 b among others. This stands in contrast to the work of Iro & Deming (2010) who, with another simple radiative model, found little ringing in their synthetic light curves of these three planets. Kataria et al. (2013) examined the phenomenon in a theoretical case with a 3D GCM coupled to radiative transfer, highlighting the importance of the viewing geometry. Ringing has not yet been observed on any exoplanet though. de Wit et al. (2016) and Lewis et al. (2017) used the lack of ringing in Spitzer observations of HD 80606 b to constrain its rotation period and infer the likely presence of clouds, demonstrating that even the non-detection of ringing holds valuable information in some systems. As these Spitzer observations had a short baseline up to two days after the periastron passage and shorter still at the longer wavelengths more suitable to observe ringing, this does not exclude the possibility that a future JWST observation with a longer baseline may detect ringing.
@@
No recent work employing a 3D GCM with full coupling to a non-grey radiative transfer implementation has simulated multiple known eccentric exoplanets to predict their photometric signals. Therefore, it is as of yet unclear how many and which planets are likely to display ringing. In this study, we generate synthetic light curves of five highly eccentric transiting exoplanets to examine whether JWST would be able to detect any ringing. Our ultimate goal is to inform the planning of JWST observing time by identifying viable targets and determining the baseline required for a ringing observation.
@@ 2 Eccentric Exoplanet GCM
We employ the atmospheric model expeRT/MITgcm (Carone et al., 2020; Schneider et al., 2022b), featuring 3D hydrodynamics coupled to a full non-grey radiative transfer solution using the pre-calculated opacity tables. It is built upon the MITgcm dynamical core of Adcroft et al. (2004) which solves the primitive equations of hydrodynamics assuming an ideal gas. The model also assumes the planet to be a rigidly rotating sphere. We use a C32 cubed-sphere grid with 47 vertical layers, the upper 41 of which are logarithmically spaced in pressure between and 100 bar and the lower six linearly spaced in steps of 100 bar down to 700 bar. A dynamical time-step of 25 s reliably obeys the Courant-Friedrichs-Lewy stability criterion (Courant et al., 1928). To smoothen small scale noise, a fourth-order Shapiro filter with a 1 day time-scale is applied (Shapiro, 1971). At the top boundary, unphysical gravity waves are dampened by a soft sponge layer, as described by Schneider et al. (2022b). Meanwhile, we apply Rayleigh friction below 200 bar to emulate magnetic drag, which Carone et al. (2020) found improves the stability of the model. We also employ the heating term due to this parameterised drag introduced by Schneider et al. (2022a).
@@
expeRT/MITgcm features a radiative transfer treatment including scattering using the Feautrier method with accelerated lambda iteration. Following Schneider et al. (2022b), we use a radiative time-step of 100 s, meaning the radiation field is only updated after every fourth dynamical time-step. Moreover, the radiative flux is only computed for every second column and then interpolated in between. We use six correlated-k bins (Goody et al., 1989) with sixteen Gaussian quadrature points each to discretise our 0.26 – 300 m wavelength range (Schneider et al., 2022b). We use the pre-computed opacity table of the chemical equilibrium model of Mollière et al. (2019), which minimises the Gibbs free energy on a grid of 1000 temperatures at our 47 pressures. It includes the molecules HO, CO, CH, NH, CO, HS, HCN, PH and FeH and the atoms Na and K as line opacity sources, Rayleigh scattering by H and He and collisionally induced absorption by interactions between H-H, H-He and H. TiO and VO are not added in any of these models as they often produce temperature inversions in the models (e.g. Showman et al., 2009; Lewis et al., 2014; Kataria et al., 2015; Schneider et al., 2022b) which are only observed on the hottest exoplanets (Heng & Marley, 2018; Arcangeli et al., 2018; Parmentier et al., 2018; Lothringer et al., 2018). The effects of clouds and hazes are not included in our model. Since measurements of the obliquity are available for few exoplanets and typically have a large uncertainty (Rauscher, 2017; Ohno & Zhang, 2019b; Adams et al., 2019; Bryan et al., 2020, 2021), we neglect obliquity in our model. Finally, we assume each atmosphere has a solar chemical composition, as is commonly done (e.g. Showman et al., 2009; Kataria et al., 2013; Amundsen et al., 2016; Lewis et al., 2017).
@@ 2.2 Implementation of eccentricity and non-synchronous rotation
We extended expeRT/MITgcm to allow for non-synchronous rotation and eccentric orbits as it was limited to tidally locked planets before. On tidally locked planets, the incoming irradiation is constant in intensity and always centred on the same point. On an eccentric and non-synchronously rotating planet on the other hand, the irradiation varies in two ways. Firstly, the intensity changes as with distance which depends on the true anomaly as . Herein, is the semi-major axis and the eccentricity. Secondly, the position of the substellar point in the planet’s co-rotating reference frame as with the time and the rotation period.
Since there is no analytical formula to calculate for any given , we apply an iterative Newton-Raphson method to the Kepler equation (Gregory, 2006) until the error in time is smaller than of the orbital period at each radiative time-step. This method needs no more than ten iterations, making it both precise and computationally cheap. However, the radiative transfer requires more iterations when the temperature structure changes due to the changing irradiation, which can increase the total computation time of the model by a factor of up to two.
@@ 2.3 Setup of the simulations
In this study, we examine five exoplanets with a high eccentricity covering a range of orbital periods, including a number of targets that have been studied before and are to be revisited by JWST. The chosen planets are HAT-P-2 b (Bakos et al., 2007), HD 80606 b (Naef et al., 2001), TOI-3362 b (Dong et al., 2021), TOI-4127 b (Gupta et al., 2023) and HD 17156 b (Fischer et al., 2007). The relevant planetary, orbital and stellar parameters are listed in Table 1. The argument of periastron defines the viewing geometry, as visualised in Fig. 1. The rotation period of the planets are computed using the commonly used pseudo-synchronous period of Hut (1981) (e.g. Kataria et al., 2013; Lewis et al., 2014, 2017; Tsai et al., 2023b) given in Eq. 1, where the rotation period is synchronised with the orbital motion around the periastron passage since that is where the tidal effects are strongest. As a result, the substellar point moves little during the periastron passage.
@@
Upon initialisation, the wind speed is zero everywhere. The temperature is set to be horizontally homogeneous with a vertical profile comprised of three parts. The upper atmosphere at pressures smaller than 1 bar is given constant temperature based on the analytical model of Parmentier et al. (2015). That model requires an estimate of the intrinsic temperature, which is computed using the relation developed in Thorngren et al. (2019). The layers deeper than 10 bar follow an adiabatic profile with a temperature of 1400 K at 1 bar, which Sainsbury-Martinez et al. (2019) found to be appropriate for the tidally locked hot Jupiter HD 209458 b. For the exoplanets modelled in this study, which have lower mean equilibrium temperatures than HD 209458 b, this deep temperature profile is likely too hot. However, Sainsbury-Martinez et al. (2019) advise using too hot a profile as atmospheric models converge more rapidly when cooling down than when heating up. Between 1 and 10 bar, the temperature profile is linearly interpolated in pressure.
@@
To reduce the total computation time, each planet is first treated as if it were tidally locked at a distance of the semi-major axis for 1000 d to let the temperature structure partly converge. Next, it is simulated for another 1000 d as an eccentric and non-synchronously rotating planet. On HD 80606 b, TOI-3362 b and TOI-4127 b, the temperature structure or jet speed at the apastron still changed considerably between consecutive orbits, so the run continued for another 1000 d, after which the jet speed or the upper atmosphere’s temperature did not monotonically change between orbits. However, the deep atmosphere’s temperature is not fully converged in any model. Schneider et al. (2022a) simulated the ultra hot Jupiter WASP-76 b for 86000 d and even after such a long run, the deep atmosphere’s temperature had not fully converged. As such, achieving complete convergence is unfeasible. Mayne et al. (2017); Carone et al. (2020) show that the dynamics of the deep atmosphere can considerably affect the upper atmosphere. However, Tsai et al. (2023b) indicate that additional internal heat from the overly hot deep atmosphere is unlikely to strongly affect the temperature and wind of the upper atmosphere of HD 80606 b, resulting in only a minor increase in planetary emission. Since the jet speed of the deep atmosphere has converged, we do not expect the incomplete temperature convergence of the deep atmosphere to notably affect our synthetic light curves. Therefore, our models are sufficiently converged for our purposes.

### v2

@@ Abstract
Context. Eccentric exoplanets offer an opportunity to study the response of an atmosphere to changing thermal forcing and the robustness of the super-rotating equatorial jet seen on tidally locked hot Jupiters. However, the atmospheric dynamics on eccentric planets strongly depend on the planetary rotation period, which is difficult to constrain observationally. The ringing phenomenon, whereby the observed emission increases and decreases after the periastron passage as the flash-heated hemisphere rotates into and out of view, can provide a tight constraint on rotation.
Aims. We studied five highly eccentric transiting exoplanets HAT-P-2 b, HD 80606 b, TOI-3362 b, TOI-4127 b and HD 17156 b to find which displays strong ringing signals that are sufficiently strong for the James Webb Space Telescope (JWST) to detect.
Methods. We implemented the treatment of eccentricity and non-synchronous rotation in the non-grey climate model expeRT/MITgcm and generated synthetic light curves.
Results. We find four detectable ringing peaks on HD 80606 b and some undetectable ringing on TOI-4127 b and HD 17156 b. The lack of clouds, photo-chemistry and obliquity in our models may have led us to overestimate the amplitude of the ringing however. The strength of the ringing signal is mostly determined by the eccentricity, planetary rotation period, planet-to-star radius ratio and apparent magnitude of the system. We searched for more exoplanets that could show ringing but found no candidates as promising as HD 80606 b.
Conclusions. We recommend prioritising HD 80606 b as a target for ringing with JWST. A baseline of five days after the periastron passage would capture three ringing peaks, which is sufficient to tightly constrain the planetary rotation period. An extension to seven days would add a fourth peak, which would allow us to verify the rotation period.
@@
In the past few decades there have been great advances in the study of exoplanetary atmospheres. The observations of the recently launched James Webb Space Telescope (JWST) have already contributed greatly to these efforts (e.g. JWST Transiting Exoplanet Community Early Release Science Team et al., 2023; Feinstein et al., 2023; Tsai et al., 2023a; Dyrek et al., 2024). Because of the complicated nature of exoplanetary atmospheres, observations should be compared to global circulation models (GCMs) to be interpreted properly. These GCMs require well-constrained orbital and planetary parameters, such as the planet’s radius, rotation period, and orbital eccentricity, as input (Showman et al., 2020). If the planet transits its parent star, most of these quantities can be reliably measured, with the notable exception of the rotation period (Sing, 2018).
@@
Most studies so far have focussed on hot Jupiters because their size and brightness makes it relatively easy to achieve a high observational signal-to-noise (Showman et al., 2020). As a result of their close-in orbits, hot Jupiters are often on a tidally circularised orbit. Based on timescale arguments, it can then safely be assumed that the planet’s rotation period is equal to the orbital period (Hut, 1981; Socrates et al., 2012; Rauscher, 2017) and that its obliquity is zero (Rauscher, 2017; Rauscher et al., 2023). Consequently, the irradiation these planets receive is constant in intensity and always illuminates the same hemisphere. On the dayside hemisphere, the illumination creates a pattern of geostrophic winds that transport angular momentum towards the equator (Showman & Polvani, 2010, 2011). Consequently, these winds form a remarkable circulation pattern that is dominated by an eastward flow of gas centred on the equator and extending around the planet, which typically reaches speeds of several km s-1. This flow is called the equatorial super-rotating jet. The existence of such jets is well established both theoretically (e.g. Showman & Guillot, 2002; Showman et al., 2009; Tsai et al., 2014; Amundsen et al., 2016; Carone et al., 2020; Showman et al., 2020) and observationally (e.g. Knutson et al., 2007; Snellen et al., 2010; Zellem et al., 2014; Louden & Wheatley, 2015; May et al., 2022). Because of the major role of the Coriolis force in atmospheric dynamics, the planet’s rotation period is a decisive parameter in the shaping of these circulation patterns (e.g. Showman et al., 2009; Showman & Polvani, 2010; Rauscher & Kempton, 2014; Showman et al., 2015; Baeyens et al., 2021; May et al., 2022).
@@
Planets on eccentric orbits can no longer be assumed to rotate synchronously (Hut, 1981). Despite the increased difficulty, studying eccentric and non-synchronously rotating exoplanets is worthwhile as it allows one to examine the response of the atmosphere’s dynamics, chemistry, and cloudiness to changing thermal forcing (Kataria et al., 2013; Lewis et al., 2017; Tsai et al., 2023b). Moreover, transit spectroscopy of cool planets on far-out, circular orbits is rarely available as the distance to the star reduces the likelihood that a transit would take place. For an eccentric planet, on the other hand, a transit is more likely to take place around periastron, when the atmosphere still contains traces of the cooler regime around apastron that is otherwise difficult to observe. Finally, about half of the known Jupiter-like planets are eccentric, especially those with a longer orbital period due to the distance dependence of tidal circularisation (Malmberg & Davies, 2009; Dawson & Murray-Clay, 2013; Bitsch et al., 2020).
@@
A number of previous studies have examined the role of eccentricity and non-synchronous rotation in the atmospheric circulation of hot and warm Jupiters. Showman et al. (2009), Rauscher & Kempton (2014), and Showman et al. (2015) used 3D GCMs to study the effect of non-synchronous rotation, concluding that a shorter rotation period weakens the equatorial jet. They disagree on the effect of an increased rotation period, though, with Rauscher & Kempton (2014) claiming that a westward flow develops and Showman et al. (2009, 2015) finding that it leads to significantly broader and faster eastward jets. Nonetheless, these studies all show that the rotation period plays a major role in atmospheric circulation. In a similar vein, Kataria et al. (2013) examined the impact of eccentricity in a theoretical scenario, pointing out that while jets are still ubiquitous, they tend to be faster the greater the eccentricity is and change quite notably in the upper atmosphere throughout the orbit. This is in agreement with the results of Lewis et al. (2014), who simulated the eccentric planet HAT-P-2 b, and Lewis et al. (2017) and Tsai et al. (2023b), who examined the extremely eccentric HD 80606 b. The latter two studies show that the jet speed on HD 80606 b decreases between periastron passages. Lewis et al. (2017) also highlight the impact of the rotation period on the circulation, composition, and lofting of clouds and, consequently, the photometric observations of highly eccentric planets.
@@
During the periastron passage, when an eccentric planet is briefly but intensely heated, most of the flash-heating occurs on the side facing the star. As the planet moves farther from the star and cools, that heated hemisphere rotates into and out of view of an observer on Earth. Consequently, the total brightness of the star and planet combined may display a number of local maxima, which is known as ringing. Thus, by studying the system’s light curve after the periastron passage, one can constrain the planet’s rotation period as well as how efficiently the temperature is homogenised by circulation. Using a simple thermal model without consistent dynamics or radiative transfer, Cowan & Agol (2011) predicted the presence of such ringing on the exoplanets HAT-P-2 b, HD 80606 b, and HD 17156 b, among others. This stands in contrast to the work of Iro & Deming (2010), who, with another simple radiative model, found little ringing in their synthetic light curves of these three planets. Kataria et al. (2013) examined the phenomenon in a theoretical case with a 3D GCM coupled to radiative transfer, highlighting the importance of the viewing geometry. Ringing has not yet been observed on any exoplanet, though. de Wit et al. (2016) and Lewis et al. (2017) used the lack of ringing in Spitzer observations of HD 80606 b to constrain its rotation period and infer the likely presence of clouds, demonstrating that even the non-detection of ringing holds valuable information in some systems. As these Spitzer observations had a short baseline – a maximum of two days after the periastron passage and shorter still at the longer wavelengths more suitable to observe ringing – this does not exclude the possibility that a future JWST observation with a longer baseline may detect ringing.
@@
No recent work employing a 3D GCM with full coupling to a non-grey radiative transfer implementation has simulated multiple known eccentric exoplanets to predict their photometric signals. Therefore, it is as of yet unclear how many and which planets are likely to display ringing. In this study we generated synthetic light curves of five highly eccentric transiting exoplanets to examine whether JWST would be able to detect any ringing. Our ultimate goal is to inform the planning of JWST observing time by identifying viable targets and determining the baseline required for a ringing observation.
@@ 2 Eccentric exoplanet GCM
We employed the atmospheric model expeRT/MITgcm (Carone et al., 2020; Schneider et al., 2022b), which features 3D hydrodynamics coupled to a full non-grey radiative transfer solution. It is built upon the MITgcm dynamical core of Adcroft et al. (2004), which solves the primitive equations of hydrodynamics assuming an ideal gas. The model also assumes the planet to be a rigidly rotating sphere. We used a C32 cubed-sphere grid with 47 vertical layers, the upper 41 of which are logarithmically spaced in pressure between and 100 bar and the lower six linearly spaced in steps of 100 bar down to 700 bar. A dynamical time-step of 25 s reliably obeys the Courant-Friedrichs-Lewy stability criterion (Courant et al., 1928). To smoothen small-scale noise, a fourth-order Shapiro filter with a 1 day timescale is applied (Shapiro, 1971). At the top boundary, unphysical gravity waves are dampened by a soft sponge layer, as described by Schneider et al. (2022b). Meanwhile, we applied Rayleigh friction below 200 bar to emulate magnetic drag, which Carone et al. (2020) find improves the stability of the model. We also employed the heating term due to this parameterised drag introduced by Schneider et al. (2022a).
@@
expeRT/MITgcm features a radiative transfer treatment including scattering using the Feautrier method with accelerated lambda iteration. Following Schneider et al. (2022b), we used a radiative time-step of 100 s, meaning the radiation field is only updated after every fourth dynamical time-step. Moreover, the radiative flux is only computed for every second column and then interpolated in between. We used six correlated-k bins (Goody et al., 1989) with sixteen Gaussian quadrature points each to discretise our 0.26 – 300 m wavelength range (Schneider et al., 2022b). We employed the pre-computed opacity table of the chemical equilibrium model of Mollière et al. (2019), which minimises the Gibbs free energy on a grid of 1000 temperatures at our 47 pressures. It includes: the molecules H2O, CO2, CH4, NH3, CO, H2S, HCN, PH3, and FeH and the atoms Na and K as line opacity sources; Rayleigh scattering by H2 and He; and absorption collisionally induced by interactions between H2-H2, H2-He, and H-. TiO and VO were not added in our models as they often produce temperature inversions in the models (e.g. Showman et al., 2009; Lewis et al., 2014; Kataria et al., 2015; Schneider et al., 2022b) and such inversions are only observed on the hottest exoplanets (Heng & Marley, 2018; Arcangeli et al., 2018; Parmentier et al., 2018; Lothringer et al., 2018). The effects of clouds and hazes were not included in our model. Since measurements of the obliquity are available for few exoplanets and typically have a large uncertainty (Rauscher, 2017; Ohno & Zhang, 2019b; Adams et al., 2019; Bryan et al., 2020, 2021), we neglected obliquity in our model. Finally, we assumed each atmosphere has a solar chemical composition, as is commonly done (e.g. Showman et al., 2009; Kataria et al., 2013; Amundsen et al., 2016; Lewis et al., 2017).
@@ 2.2 Implementation of eccentricity and non-synchronous rotation
We extended expeRT/MITgcm to allow for non-synchronous rotation and eccentric orbits as it was limited to tidally locked planets before. On tidally locked planets, the incoming irradiation is constant in intensity and always centred on the same point. On an eccentric and non-synchronously rotating planet on the other hand, the irradiation varies in two ways. Firstly, the intensity changes as , with distance , which depends on the true anomaly as . Herein, is the semi-major axis and the eccentricity. Secondly, the position of the substellar point in the planet’s co-rotating reference frame as with the time and the rotation period.
Since there is no analytical formula to calculate for any given , we applied an iterative Newton-Raphson method to the Kepler equation (Gregory, 2006) until the error in time was smaller than of the orbital period at each radiative time-step. This method needs no more than ten iterations, making it both precise and computationally cheap. However, the radiative transfer requires more iterations when the temperature structure changes due to the changing irradiation, which can increase the total computation time of the model by a factor of up to two.
@@ 2.3 Setup of the simulations
In this study we examined five exoplanets with a high eccentricity covering a range of orbital periods, including a number of targets that have been studied before and are to be revisited by JWST. The chosen planets are HAT-P-2 b (Bakos et al., 2007), HD 80606 b (Naef et al., 2001), TOI-3362 b (Dong et al., 2021), TOI-4127 b (Gupta et al., 2023), and HD 17156 b (Fischer et al., 2007). The relevant planetary, orbital and stellar parameters are listed in Table 1. The argument of periastron defines the viewing geometry, as visualised in Fig. 1. The rotation periods of the planets were computed using the commonly used pseudo-synchronous period of Hut (1981) (e.g. Kataria et al., 2013; Lewis et al., 2014, 2017; Tsai et al., 2023b) given in Eq. 1, where the rotation period is synchronised with the orbital motion around the periastron passage since that is where the tidal effects are strongest. As a result, the substellar point moves little during the periastron passage:
@@
Upon initialisation, the wind speed was zero everywhere. The temperature was horizontally homogeneous with a vertical profile comprised of three parts. The upper atmosphere at pressures smaller than 1 bar was given a constant temperature based on the analytical model of Parmentier et al. (2015). That model requires an estimate of the intrinsic temperature, which is computed using the relation developed in Thorngren et al. (2019). The layers deeper than 10 bar followed an adiabatic profile with a temperature of 1400 K at 1 bar, which Sainsbury-Martinez et al. (2019) find to be appropriate for the tidally locked hot Jupiter HD 209458 b. For the exoplanets modelled in this study, which have lower mean equilibrium temperatures than HD 209458 b, this deep temperature profile is likely too hot. However, Sainsbury-Martinez et al. (2019) advise using too hot a profile as atmospheric models converge more rapidly when cooling down than when heating up. Between 1 and 10 bar, the temperature profile is linearly interpolated in pressure.
@@
To reduce the total computation time, each planet was first treated as if it were tidally locked at a distance of the semi-major axis for 1000 d to let the temperature structure partly converge. Next, it was simulated for another 1000 d as an eccentric and non-synchronously rotating planet. On HD 80606 b, TOI-3362 b, and TOI-4127 b, the temperature structure or jet speed at the apastron still changed considerably between consecutive orbits, so the run continued for another 1000 d, after which the jet speed or the upper atmosphere’s temperature did not monotonically change between orbits. However, the deep atmosphere’s temperature was still not fully converged in any model. Schneider et al. (2022a) simulated the ultra hot Jupiter WASP-76 b for 86000 d and even after such a long run, the deep atmosphere’s temperature had not fully converged. As such, achieving complete convergence is unfeasible. Mayne et al. (2017) and Carone et al. (2020) show that the dynamics of the deep atmosphere can considerably affect the upper atmosphere. However, Tsai et al. (2023b) indicate that additional internal heat from the overly hot deep atmosphere is unlikely to strongly affect the temperature and wind of the upper atmosphere of HD 80606 b, resulting in only a minor increase in planetary emission. Since the jet speed of the deep atmosphere has converged, we do not expect the incomplete temperature convergence of the deep atmosphere to notably affect our synthetic light curves. Therefore, our models are sufficiently converged for our purposes.